an airplane flies 750 miles with the wind in the same amount of time the airplane could fly 625 mi against the wind. assuming the wind speed is constant, and given that the plane can fly 275 mph in still air, find the speed of the wind

Answer by richwmiller(17219)

r*t=d
(275+w)*t=750
t=750/(275+w)
(275-w)*t=625
t=625/(275-w)
t=750/(275+w)
t=625/(275-w)
750/(275+w)=625/(275-w)
w = 25